ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 62G07, 65C10, 11K45 Herrmann, Eva, Technische Universität Darmstadt, Germany Visualization of empirical test results for pseudorandom number generators. Besides theoretical analysis, empirical tests are important tools for the assessment of pseudorandom number generators. These tests often produce a large amount of numerical results. Hence, it may become quite difficult to extract important information. Here, we will show how nonparametric curve estimation methods can help to detect deficiencies of pseudorandom number generators by giving a detailed view on the distribution of some test statistics. Part of these results is joint work with S. Wegenkittl and J. Eichenauer-Herrmann.